OpenStudy (anonymous):

Match each equation with the ordered pair that is a solution. Match Term Definition y equals x plus four A) six comma negative three y equals x minus nine B) three comma five y equals two x minus one C) negative two comma two y equals four x plus three D) negative one comma negative one

hero (hero):

y = x + 4 y = x - 9 y = 2x + 1 y = 4x + 3 A. (6,-3) B. (3,5) C. (-2,2) D. (-1,-1) To match test each one of the points with each equation. Any test that results in a true condition is a match.

hero (hero):

For example, test the point \((\color\green{6},\color\red{3})\) with the equation \(\color\red{y}= \color\green{x} + 4\) .... \(\color\red{3} = \color\green{6} + 4\) \(3 = 10\) False

hero (hero):

So obviously the point (6,3) is not in y = x + 4. Test the other points to find out which one matches.

OpenStudy (anonymous):

y = x + 4 Answer is C..

hero (hero):

Test y = x + 4 at point (-2,2) 2 = -2 + 4 2 = 4 - 2 2 = 2 True

hero (hero):

Okay, now follow that same testing process for other points and equations until you've matched them all correctly :)

OpenStudy (anonymous):

okay, Thank you.